OpenStudy (anonymous):

how do you turn sqrt(3)/3 into 1/sqrt(3)? I'm in Calculus, but I've forgotten how to do something simple like this.

OpenStudy (zarkon):

\[\frac{\sqrt{3}}{3}=\frac{\sqrt{3}}{\sqrt{3}\sqrt{3}}=\frac{1}{\sqrt{3}}\]

OpenStudy (zarkon):

\[\sqrt{3}{\sqrt{3}}=\sqrt{3}^2=3\]

OpenStudy (anonymous):

Ah, okay. That makes sense. \[\frac{ \sqrt{3} }{ 3}\] is the answer after you rationalize the denominator, right?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{3} }\] is the unrationalized answer.